OpenStudy (jenniferjuice):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (jenniferjuice):

y = -2x + 5 y = 2x - 3 y = 2x + 6 y = -2x + 2

OpenStudy (jenniferjuice):

@tcarroll010 @thomaster @tyteen4a03 @turtlemaroon @Tuiti @tayking can you guys help me please ? :)

OpenStudy (jenniferjuice):

? anyone ?

OpenStudy (jenniferjuice):

is it y = -2x + 2 ?

OpenStudy (anonymous):

first you have to find the slope slope formula is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

then I believe you use point slope form y-y1=m(x-x1)

OpenStudy (jenniferjuice):

so ... ?

OpenStudy (anonymous):

I got -2 for the slope

OpenStudy (jenniferjuice):

so is it y = -2x + 5 ?

OpenStudy (jenniferjuice):

?

OpenStudy (anonymous):

i think it is y=2x+6 http://openstudy.com/study#/updates/5080287ce4b0b8b0cacd8794